can you please help me solve this equation, (2-x)(x+4)≥0
Answer by Cromlix(4381)

Hi there,
(2 - x)(x + 4) = 8 - 2x - x^2
Curve cuts x axis at:-
2 - x = 0
-x = -2
x = 2
x + 4 = 0
x = -4
-4 <= x =< 2
This is the part of the curve on and above the x axis (x => 0)
Hope this helps :-)